OpenStudy (anonymous):

simplify 1. x^2-5x+4/x-1 2. 9x^2-3x/3x i got 3x-1 for this 3. x-y/x+y+x+y/x-y i got 2*x is this correct

OpenStudy (anonymous):

your question doesnt make sence??? lol

OpenStudy (mertsj):

1. \[\frac{x^2-5x+4}{x-1}=\frac{(x-4)(x-1)}{x-1}=x-4\]

OpenStudy (anonymous):

lol i wrote it how it was the second two i attempted myself

OpenStudy (anonymous):

like how Mertsj did

OpenStudy (mertsj):

2. Correct 3. \[\frac{x-y}{x+y}+\frac{x+y}{x-y}=\frac{(x-y)^2+(x+y)^2}{(x-y)(x+y)}=\] \[\frac{x^2-2xy+y^2+x^2+2xy+y^2}{(x-y)(x+y)}=\frac{2x^2+2y^2}{x^2-y^2}\]

OpenStudy (anonymous):

thanks alot !!! :)

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

:)